ABC Acquired 80% of XYZ shares on 1 January 2019. The purchase consideration was Sh. 11 million cash upfront and an additional Sh. 800,000 payable if XYZ maintains profit margin of 30% for the period ending 31 December 2019. XYZ has one million shares and was trading at Sh. 14 per share prior to acquisition. The fair value of net identifiable assets of XYZ on acquisition date was Sh. 12 million. Required: Calculate Goodwill recognized on 1 January 2019 using the fair value method.


1
Expert's answer
2021-05-17T16:26:07-0400

"Goodwill = investment amount — the sum of the net assets of the subsidiary\times the percentage of ownership."

investment amount =11+0.8=11.8

"Goodwill = investment amount — the sum of the net assets of the subsidiary\times the percentage of ownership.=11.8-12\times0.8=2.2"